The sum of two even numbers is even.
Using Direct proof method
1
Expert's answer
2020-11-10T20:00:57-0500

Let a and b be even numbers.

a = 2*c and b = 2*d for some integers c and d.

a+b = 2*c+2*d = 2*(c+d) = 2*k (where k is also some integer)

So, a+b is even